Determine graphically the magnitude and direction of the
resultant of the coplanar forces given which are acting at a
point.
Q3: Force A, 35N acting horizontally to the right, force B,
15N acting at 160° to force A; force C, 18N acting at 280°
to force A.
Q4: Force 1, 23 kN acting at 60° to the horizontal, force 2,
30 kN acting at 32° to force 1, force 3, 25 kN acting at 70°
to force 2.
1
Expert's answer
2019-12-16T11:00:54-0500

Q3

"result vector =27 N"

"\alpha=332^o"

"resultvector=70 kN"

"\alpha=95^o"